#6 a. Margaret’s regular hourly wage rate is $20, and she receives an hourly rate of...

#6

a. Margaret’s regular hourly wage rate is $20, and she receives an hourly rate of $30 for work in excess of 40 hours. During a January pay period, Margaret works 49 hours. Margaret’s federal income tax withholding is $97, and she has no voluntary deductions.

Prepare the employer’s journal entry to record payroll taxes for the period. Ignore unemployment taxes. (Credit account titles are automatically indented when amount is entered. Do not indent manually. Assume FICA rate of 7.65%. Round the answers to 2 decimal places, e.g. 15.25.)

b. Blossom Company issues $410,000, 10-year, 7% bonds at 95.

Prepare the journal entry to record the sale of these bonds on March 1, 2019. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

c. Pina Colada Corp. issues $330,000, 20-year, 6% bonds at 104.

Prepare the journal entry to record the sale of these bonds on June 1, 2019. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1
a
Gross pay 1070 =(40*20)+(49-40)*30
Journal Debit Credit
Payroll Taxes Expense 81.86 =1070*7.65%
        FICA Taxes Payable 81.86
b
March 1, 2019
Journal Debit Credit
Cash 389500 =410000*0.95
Discount on Bonds payable 20500
       Bonds payable 410000
c
June 1, 2019
Journal Debit Credit
Cash 343200 =330000*1.04
       Bonds payable 330000
       Premium on Bonds payable 13200